OpenStudy (anonymous):

solve the following equation for real values of x by completing the square. the equations are: x^2-3x+28 and 9x^2-12x+5=0

OpenStudy (anonymous):

really? it is easiest to use formula if "middle term" is odd, but you can still complete the square if you like

OpenStudy (anonymous):

i probably sound really clueless, but how do i complete a square?

OpenStudy (anonymous):

like this:

OpenStudy (anonymous):

i assume the first one is \[x^2-3x+28=0\] yes?

OpenStudy (anonymous):

it will have no real solutions. are you allowing complex answers?

OpenStudy (anonymous):

yes, any answers are allowed

OpenStudy (anonymous):

well actually only real numbers are acceptable. if that makes any sense...

OpenStudy (anonymous):

ok so i would use quadratic formula. but if you want to complete the square start with \[x^2-3x=-28\] \[(x-\frac{3}{2})^2=-28+(\frac{3}{2})^2\] \[(x-\frac{3}{2})^2=-\frac{103}{4}\]

OpenStudy (anonymous):

in that case stop right here because the left hand side is a perfect square and hence not negative, right hand side is negative. no solution

OpenStudy (anonymous):

okay i see, thanks so much!

OpenStudy (anonymous):

the next one you have to divide everything by 9 as a first step.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

okay i see.